### v2

# Maximal $d$-spectra via Priestley duality

Abstract.
We use Priestley duality as a new tool to study maximal $d$-spectra of arithmetic frames, both with and without units. We pay special attention to when the maximal $d$-spectrum is compact or Hausdorff. Various necessary and sufficient conditions are given, including a construction of an arithmetic frame with a unit whose maximal $d$-spectrum is not Hausdorff, thus resolving an open problem in the literature.
Key words and phrases:
Pointfree topology, Priestley duality, arithmetic frame, $d$-nucleus, $d$-ideal
2020 Mathematics Subject Classification:
18F70; 06D22; 06E15; 06F20; 54D10; 54D30;
1. Introduction
The space of maximal $d$-ideals of an archimedean Riesz space with a weak order unit, equipped with the hull-kernel topology, has been well studied and is known to be a compact Hausdorff space (see, e.g., [22]). Motivated by this, Martinez and Zenk [29] initiated the study of $d$-elements in an arbitrary arithmetic frame. These elements, denoted $dL$ or $L_{d}$, form a sublocale of the arithmetic frame $L$. The corresponding nucleus was coined the $d$-nucleus. The frame $L_{d}$ and the spectrum $\max L_{d}$ of maximal $d$-elements were further studied by various authors (see, e.g., [15, 16, 8, 9]). It is known that $\max L_{d}$ is a compact $T_{1}$-space provided $L$ has a unit. In [8], it was left open whether $\max L_{d}$ is Hausdorff. Although some characterizations of the Hausdorff separation for $\max L_{d}$ were recently established in [9], the question remained open. Our aim is to answer it in the negative.
Our main tool is Priestley duality for frames. Priestley originally developed her duality for bounded distributive lattices [33, 34]. It was restricted to the category of frames by Pultr and Sichler [36], and later the Priestley spaces of arithmetic frames were characterized in [13]. Building on this work, we describe the subset $N_{d}$ of the Priestley space $X$ of an arithmetic frame $L$ corresponding to the $d$-nucleus on $L$. We also describe the subset $Y_{d}$ of $N_{d}$ corresponding to the spectrum $pt(L_{d})$ of points of the sublocale $L_{d}$. We show that the minimum of $Y_{d}$ is in a one-to-one correspondence with the maximal $d$-elements of $L$, thus yielding a homeomorphism between $\min Y_{d}$ and $\max L_{d}$. This allows us to study $\max L_{d}$ in the language of Priestley spaces. Our main results include the description of the nucleus on $L$ whose fixpoints are the frame of opens of $\min Y_{d}$, the characterization of the soberification of $\min Y_{d}$, and the construction of an arithmetic frame $L$ with a unit such that $\min Y_{d}$ is not Hausdorff. This yields that $\max L_{d}$ is not Hausdorff, thus resolving the open question mentioned above. We also give a necessary and sufficient condition for $\max L_{d}$ to be Hausdorff. This characterization remains valid even if $L$ doesn’t have a unit, provided $\max L_{d}$ is locally compact. We also investigate the compactness of $\min Y_{d}$, and hence of $\max L_{d}$, in comparison with the existence of a unit. Our approach raises new open questions and highlights the need for further study of maximal $d$-spectra of arithmetic frames (see the end of the paper).
The paper is structured as follows. In Section 2, we recall Priestley duality for arithmetic frames, along with useful definitions and results from the literature. In Section 3, we revisit the relationship between nuclei and sublocales, as well as their description in the language of Priestley spaces. In particular, we show how to use Priestley duality to give alternative proofs of two existing results in the literature; Johonstone’s lemma that each Scott open filter arises as the dense elements of a nucleus and the Isbell Density Theorem.
In Section 4, we characterize inductive nuclei in the language of Priestley spaces, as well as provide the dual description of the $d$-nucleus on an arithmetic frame $L$. In Section 5, we study the maximum of the localic part of the Priestley dual of $L$, which yields a new characterization of when the sublocale $L_{d}$ is regular. In Section 6, we delve into the investigation of the spectrum $\max L_{d}$ of maximal $d$-elements of an arithmetic frame. We establish a homeomorphism between $\max L_{d}$ and $\min Y_{d}$, thus giving us a new tool to study the maximal $d$-spectrum of $L$. We show that the frame of open sets of $\min Y_{d}$ can be realized as a sublocale of $L$ and describe the corresponding nuclear subset of $X$. In addition, we prove that the localic part of this nuclear subset is the soberification of $\min Y_{d}$.
In Section 7, we study the topological properties of $\min Y_{d}$ in comparison to what is known about $\max L_{d}$. We describe compact subsets of $\min Y_{d}$, which allows us to characterize when an arithmetic frame has a unit using Priestley duality. This in particular yields that in the presence of a unit, $\min Y_{d}$ is a compact space.
Finally, in Section 8, we explore the Hausdorff separation for the space $\min Y_{d}$. An example of an arithmetic frame $L$ with a unit is constructed such that $\max L_{d}$ and hence $\min Y_{d}$ is not Hausdorff, thus resolving an open question from [8]. Furthermore, we give a characterization of when $\min Y_{d}$ is Hausdorff, which generalizes to arithmetic frames without units, provided $\min Y_{d}$ is locally compact. The article concludes with several open questions that the authors are looking into to further the study of $\min Y_{d}$.
2. Priestley duality for frames
A frame is a complete lattice $L$ such that
$a\wedge\bigvee S=\bigvee\{a\wedge s\mid s\in S\}$
for all $a\in L$ and $S\subseteq L$. A frame homomorphism is a map between frames that preserves finite meets and arbitrary joins. Let Frm be the category of frames and frame homomorphisms. A frame $L$ is spatial if it is isomorphic to the frame of opens of a topological space (equivalently, completely prime filters separate elements of $L$). Let SFrm be the full subcategory of Frm consisting of spatial frames.
An element $a\in L$ is compact if for each $S\subseteq L$, from $a\leq\bigvee S$ it follows that $a\leq\bigvee T$ for some finite $T\subseteq S$. Let $K(L)$ be the set of compact elements of $L$. Then $L$ is algebraic provided
$a=\bigvee\{b\in K(L)\mid b\leq a\}$
for each $a\in L$. It is well known (see, e.g., [29, Rem. 3.4]) that every algebraic frame is spatial. A frame homomorphism $h:L\to M$ is coherent if $h[K(L)]\subseteq K(M)$. Let AlgFrm be the category of algebraic frames and coherent frame homomorphisms between them. An algebraic frame $L$ is arithmetic if $K(L)$ is closed under binary meets. Let AriFrm be the full subcategory of AlgFrm consisting of arithmetic frames.
A space is zero-dimensional if it has a basis of clopen sets. A Stone space is a zero-dimensional, compact, Hausdorff space. A Priestley space is a pair $(X,\leq)$ such that $X$ is a Stone space and $\leq$ is a partial order on $X$ such that the Priestley separation holds:
If $x\nleq y$ then there is a clopen upset $U$ of $X$ containing $x$ and missing $y$.
A Priestley morphism is a continuous order-preserving map between Priestley spaces. Let Pries be the category of Priestley spaces and Priestley morphisms. Let DLat be the category of bounded distributive lattice and bounded lattice homomorphisms.
Theorem 2.1 (Priestley duality).
DLat and Pries are dually equivalent.
Remark 2.2.
The functors $\mathscr{X}:{\sf DLat}\to{\sf Pries}$ and $\mathscr{D}:{\sf Pries}\to{\sf DLat}$ establishing Priestley duality are described as follows.
  • The Priestley space of a bounded distributive lattice $D$ is the set $X_{D}$ of prime filters of $D$ ordered by inclusion and topologized by the basis $\{\varphi(a)\setminus\varphi(b)\mid a,b\in D\}$, where $\varphi$ is the Stone map defined by $\varphi(a)=\{x\in X_{D}\mid a\in x\}$ for all $a\in D$. The functor $\mathscr{X}$ assigns to each $D\in{\sf DLat}$ its Priestley space $X_{D}$, and to each bounded lattice homomorphism $h:D\to E$ the Priestley morphism $h^{-1}:X_{E}\to X_{D}$.
  • The functor $\mathscr{D}$ assigns to each Priestley space $X$ the bounded distributive lattice ${\sf ClopUp}(X)$ of clopen upsets of $X$ and to each Priestley morphism $f:X\to Y$ the bounded lattice homomorphism $f^{-1}:{\sf ClopUp}(Y)\to{\sf ClopUp}(X)$.
Definition 2.3.
  1. (1)
    An L-space (localic space) is a Priestley space such that the closure of each open upset is an open upset.
  2. (2)
    An L-morphism is a Priestley morphism $f:X\to Y$ between L-spaces satisfying $f^{-1}(\operatorname{cl}U)=\operatorname{cl}f^{-1}(U)$ for each open upset $U$ of $Y$.
  3. (3)
    LPries is the category of L-spaces and L-morphisms.
Remark 2.4.
Recall that frames are precisely complete Heyting algebras. By Esakia duality [19], the Priestley spaces of Heyting algebras (Esakia spaces) are those with the property that the closure of each upset is an upset. Moreover, if $X$ is the Priestley space of an Heyting algebra $L$, then $L$ is complete iff $X$ is extremally order-disconnected; that is, the closure of each open upset is open. Thus, L-spaces are precisely extremally order-disconnected Esakia spaces.
Theorem 2.5 (Pultr-Sichler duality).
Frm and LPries are dually equivalent.
Throughout we will use the following well-known facts (see, e.g., [35, 20, 2]).
Lemma 2.6.
Let $X$ be a Priestley space.
  1. (1)
    If $F\subseteq X$ is closed, then so are $\mathord{\uparrow}F$ and $\mathord{\downarrow}F$.
  2. (2)
    If $F\subseteq X$ is a closed upset, then it is an intersection of clopen upsets.
  3. (3)
    If $F\subseteq X$ is closed, then for each $x\in F$ there are $m\in\min F$ and $n\in\max F$ such that $m\leq x\leq n$, where $\min F$ and $\max F$ denote the sets of minimal and maximal points of $F$, respectively.
If in addition $X$ is an L-space, then we have:
  1. (4)
    If $F\subseteq X$ is closed, then so is $\max F$.
  2. (5)
    If $U\subseteq X$ is clopen, then so is $\mathord{\downarrow}U$.
  3. (6)
    For $\{U_{i}\}\subseteq{\sf ClopUp}(X)$,
    $\bigvee U_{i}=\operatorname{cl}\bigcup U_{i}\quad\mbox{and}\quad\bigwedge U_{i}=X\setminus\mathord{\downarrow}(X\setminus\operatorname{int}\bigcap U_{i}).$
We next describe the Priestley spaces of spatial, algebraic, and arithmetic frames.
Definition 2.7.
For an L-space $X$,
  1. (1)
    $y\in X$ is a localic point if $\mathord{\downarrow}y$ is clopen;
  2. (2)
    the set $Y$ of localic points of $X$ is the localic part of $X$;
  3. (3)
    $X$ is an SL-space (spatial L-space) if $Y$ is dense in $X$;
  4. (4)
    SLPries is the full subcategory of LPries consisting of SL-spaces.
Theorem 2.8 (see, e.g., [12, Cor. 4.10]).
SFrm and SLPries are dually equivalent.
Remark 2.9.
Let $X$ be the Priestley space of a frame $L$, and $Y$ the localic part of $X$. We topologize $Y$ by $\{U\cap Y\mid U\in{\sf ClopUp}(X)\}$. Then $Y$ is precisely the space of points of $L$ (see, e.g, [1, Prop 5.1]).
Definition 2.10.
Let $X$ be an L-space and $Y$ its localic part.
  1. (1)
    A closed upset $F\subseteq X$ is a Scott upset if $\min F\subseteq Y$.
  2. (2)
    ${\sf ClopSUp}(X)$ is the collection of clopen Scott upsets of $X$.
Scott upsets play an important role in Priestley spaces of frames as they correspond to Scott open filters on the frame side:
Theorem 2.11 ([11, Thm. 5.6]).
Let $L$ be a frame and $X$ its L-space. The poset of Scott open filters of $L$ is dually isomorphic to the poset of Scott upsets of $X$.
Definition 2.12.
Let $X$ be an L-space and $Y$ its localic part.
  1. (1)
    The core of $U\in{\sf ClopUp}(X)$ is $\operatorname{core}U=\bigcup\{V\in{\sf ClopSUp}(X)\mid V\subseteq U\}$.
  2. (2)
    $X$ is an algebraic L-space if $\operatorname{core}U$ is dense in $U$ for each $U\in{\sf ClopUp}(X)$.
  3. (3)
    An L-morphism $f:X_{1}\to X_{2}$ is coherent if $f^{-1}(\operatorname{core}U)\subseteq\operatorname{core}f^{-1}(U)$ for each $U\in{\sf ClopUp}(X_{2})$.
  4. (4)
    AlgL is the category of algebraic L-spaces and coherent L-morphisms between them.
Theorem 2.13 ([13, Thm.4.9]).
AlgFrm and AlgL are dually equivalent.
Consequently, every algebraic L-space is an SL-space.
Remark 2.14.
Let $X$ be an L-space and $U\in{\sf ClopUp}(X)$. Then $\operatorname{core}U=U$ iff $U$ is a Scott upset, which by [11, Cor. 5.4] is equivalent to $U$ being a compact element of ${\sf ClopUp}(X)$. Thus, ${\sf ClopSUp}(X)=K({\sf ClopUp}(X))$.
Definition 2.15.
  1. (1)
    An arithmetic L-space is an algebraic L-space $X$ such that
    $\operatorname{core}U\cap\operatorname{core}V=\operatorname{core}(U\cap V)$
    for all $U,V\in{\sf ClopUp}(X)$. Equivalently (see, e.g., [13, Lem. 5.2]),
    $U,V\in{\sf ClopSUp}(X)\implies U\cap V\in{\sf ClopSUp}(X).$
  2. (2)
    AriL is the full subcategory of AlgL consisting of arithmetic L-spaces.
Theorem 2.16 ([13, Thm 5.5]).
AriFrm and AriL are dually equivalent.
3. Priestley duals for nuclei and sublocales
In this section, we recall the relationship between nuclei and sublocales of a frame $L$, as well as their dual characterization as nuclear subsets of the Priestley space of $L$. Moreover, we give a dual characterization of the admissible filter of a nucleus on $L$, which yields an alternative proof of [26, Lem. 3.4(ii)]. Furthermore, we provide an alternate proof of Isbell’s Density Theorem, which utilizes Priestley duality.
Definition 3.1 (see, e.g., [25, p. 48]).
A nucleus on a frame $L$ is a map $j:L\to L$ satisfying
  1. (1)
    $a\leq ja$
  2. (2)
    $jja\leq ja$
  3. (3)
    $ja\wedge jb=j(a\wedge b)$
for all $a,b\in L$.
Let $N(L)$ be the set of nuclei on $L$. We order $N(L)$ pointwise, i.e., $j\leq k$ iff $j(a)\leq k(a)$ for all $a\in L$. With this order, it is well known that $N(L)$ is a frame (see, e.g., [25, p. 51]).
Definition 3.2 (see, e.g., [31, p. 26]).
Let $L$ be a frame. We call $S\subseteq L$ a sublocale of $L$ provided $S$ is closed under arbitrary meets and $a\to s\in S$ for all $a\in L$ and $s\in S$.
Let $S(L)$ be the set of sublocales of $L$. We order $S(L)$ by inclusion. Then $S(L)$ is dually isomorphic to $N(L)$ (see, e.g., [31, Sect. III-5]). The dual isomorphism associates with each nucleus $j$, the sublocale $S_{j}\coloneqq j[L]\in S(L)$; and with each sublocale $S$, the nucleus $j_{S}$ given by $j_{S}(a)=\bigwedge\{s\in S\mid a\leq s\}$. Nuclei, and hence sublocales, dually correspond to nuclear subsets of L-spaces introduced in [37] (see also [6, 1]):
Definition 3.3.
Let $X$ be an L-space.
  1. (1)
    A subset $N$ of $X$ is a nuclear subset provided $N$ is closed and $\mathord{\downarrow}(U\cap N)$ is clopen for each clopen subset $U$ of $X$.
  2. (2)
    Let $N(X)$ be the set of nuclear subsets of $X$ ordered by inclusion.
Theorem 3.4 ([6, Thm. 30]).
Let $L$ be a frame and $X$ its Priestley space. Then $N(L)$ is dually isomorphic to $N(X)$.
Remark 3.5.
The dual isomorphism of the previous theorem is established as follows: with each $j\in N(L)$ we associate the nuclear subset
$N_{j}\coloneqq\{x\in X\mid j^{-1}[x]=x\}\in N(X),$
and with each $N\in N(X)$ the nucleus $j_{N}\in N({\sf ClopUp}(X))$ given by $j_{N}\,U=X\setminus\mathord{\downarrow}(N\setminus U)$}. Then $j\coloneqq\varphi^{-1}\circ j_{N}\circ\varphi$ is the corresponding nucleus on $L$.
To simplify notation, we identify $N(L)$ with $N({\sf ClopUp}(X))$. Thus, each $j\in N(L)$ is identified with $j_{N_{j}}$, and hence, for $U\in{\sf ClopUp}(X)$, we have $j\,U=X\setminus\mathord{\downarrow}(N_{j}\setminus U)$.
Corollary 3.6 ([37, p. 229]).
Let $L$ be a frame and $X$ its Priestley space. Then $S(L)$ is isomorphic to $N(X)$.
Remark 3.7.
Let $L$ be a frame and $X$ its Priestley space.
  1. (1)
    If $j\in N(L)$, then $N_{j}$ seen as a subspace of $X$ is order-homeomorphic to the Priestley space of the sublocale $S_{j}$ of $L$ (see, e.g., [6, Lem. 25]).
  2. (2)
    Localic points of $X$ also known as nuclear points (see, e.g., [1, Def. 4.1]). This is because $y\in X$ is localic iff $\{y\}$ is a nuclear subset.
  3. (3)
    By [1, Lem. 4.8], the join in $N(X)$ is calculated by
    $\bigvee N_{i}=\operatorname{cl}\bigcup N_{i}$
    for $\{N_{i}\}\subseteq N(X)$.
  4. (4)
    By (3), $\operatorname{cl}Z$ is a nuclear subset of $X$ for every subset $Z\subseteq Y$ of the localic part.
Let $j$ be a nucleus on $L$. An element $a\in L$ is called $j$-dense provided $ja=1$. Let $F_{j}$ be the set of all $j$-dense elements of $L$. It is well known and straightforward to verify that $F_{j}$ is a filter of $L$.
Definition 3.8 (see, e.g., [39]).
A filter $F$ of a frame $L$ is called admissible if it is of the form $F=F_{j}$ for some $j\in N(L)$.
Remark 3.9.
In [27] admissible filters are called smooth. In [40, Thm. 25.5] it is shown that a filter is admissible iff it is free, a concept that is now known as strongly exact (see, e.g., [28]).
Let $X$ be the Priestley space of $L$. Recalling the well-known correspondence between filters of $L$ and closed upsets of $X$ (see [35, p. 54] or [4, Cor. 6.3]), let $H_{j}\coloneqq\bigcap\varphi[F_{j}]$} be the closed upset of $X$ corresponding to $F_{j}$. To describe the relationship between $N_{j}$ and $H_{j}$, we require the following lemma.
Lemma 3.10.
Let $L$ be a frame, $X$ its Priestley space, $j\in N(L)$, and $a,b\in L$.
  1. (1)
    $\varphi(a)\cap N_{j}=\varphi(ja)\cap N_{j}$.
  2. (2)
    $x\in\varphi(ja)$ iff $\mathord{\uparrow}x\cap N_{j}\subseteq\varphi(a)\cap N_{j}$.
  3. (3)
    $ja\leq jb$ iff $\varphi(a)\cap N_{j}\subseteq\varphi(b)\cap N_{j}$.
  4. (4)
    $ja=jb$ iff $\varphi(a)\cap N_{j}=\varphi(b)\cap N_{j}$.
  5. (5)
    $a$ is $j$-dense iff $N_{j}\subseteq\varphi(a)$.
Proof.
(1) The left-to-right inclusion is clear because $a\leq ja$. For the right-to-left inclusion, suppose $x\in\varphi(ja)\cap N_{j}$. Then $ja\in x$, so $a\in j^{-1}[x]$. But $j^{-1}[x]=x$ since $x\in N_{j}$. Consequently, $x\in\varphi(a)\cap N_{j}$.
(2) Let $x\in X$. We have
$x\in\varphi(ja)$ $\iff x\in X\setminus\mathord{\downarrow}(N_{j}\setminus\varphi(a))$
$\iff x\notin\mathord{\downarrow}(N_{j}\setminus\varphi(a))$
$\iff\mathord{\uparrow}x\cap(N_{j}\setminus\varphi(a))=\varnothing$
$\iff\mathord{\uparrow}x\cap N_{j}\subseteq\varphi(a)\cap N_{j}.$
(3) Suppose $ja\leq jb$. Then $\varphi(ja)\subseteq\varphi(jb)$. Consequently,
$\varphi(ja)\cap N_{j}\subseteq\varphi(jb)\cap N_{j},$
and so $\varphi(a)\cap N_{j}\subseteq\varphi(b)\cap N_{j}$ by (1). Conversely, suppose $\varphi(a)\cap N_{j}\subseteq\varphi(b)\cap N_{j}$. It suffices to show that $\varphi(ja)\subseteq\varphi(jb)$. Let $x\in\varphi(ja)$. Then $\mathord{\uparrow}x\cap N_{j}\subseteq\varphi(a)\cap N_{j}$ by (2). Therefore, $\mathord{\uparrow}x\cap N_{j}\subseteq\varphi(b)\cap N_{j}$ by assumption. Thus, $x\in\varphi(jb)$ by (2).
(4) This follows from (3).
(5) Suppose $ja=1$. Then $\varphi(ja)=X$. Therefore, by (1),
$N_{j}\cap\varphi(a)=N_{j}\cap\varphi(ja)=N_{j}.$
Thus, $N_{j}\subseteq\varphi(a)$. Conversely, suppose $N_{j}\subseteq\varphi(a)$. Then $\varphi(a)\cap N_{j}=N_{j}=\varphi(1)\cap N_{j}$. Therefore, by (4), $ja=j1=1$. ∎
Theorem 3.11.
Let $L$ be a frame, $X$ its Priestley space, and $j\in N(L)$. Then $H_{j}=\mathord{\uparrow}N_{j}$.
Proof.
Since both $H_{j}$ and $\mathord{\uparrow}N_{j}$ are closed upsets, and hence intersections of clopen upsets (see Item 2.6(2)), it is sufficient to show that $H_{j}\subseteq\varphi(a)$ iff $\mathord{\uparrow}N_{j}\subseteq\varphi(a)$ for each $a\in L$. We have
$H_{j}=\bigcap_{b\in F_{j}}\varphi(b)\subseteq\varphi(a)\iff a\in F_{j}\iff ja=1\iff N_{j}\subseteq\varphi(a)\iff\mathord{\uparrow}N_{j}\subseteq\varphi(a),$
where the first equivalence follows from compactness and the second to last equivalence from Item (5). ∎
The following result about Scott open filters was established in [26, Lem. 3.4(ii)] using transfinite induction. Since then various alternative proofs have been obtained (see, e.g., [24, Sec. 5.1] and the references therein). We utilize Priestley duality to provide a simpler proof.
Corollary 3.12.
Every Scott open filter of a frame is admissible.
Proof.
Let $L$ be a frame and $X$ its L-space. By Theorem 2.11, Scott open filters correspond to Scott upsets; and by Theorem 3.11, admissible filters correspond to closed upsets of the form $\mathord{\uparrow}N$ for some $N\in N(X)$. Therefore, it suffices to show that for each Scott upset $F$, there exists a nuclear subset $N\subseteq X$ such that $F=\mathord{\uparrow}N$. Let $N=\operatorname{cl}(F\cap Y)$. Then $N$ is nuclear by Item (4). Moreover, since $F$ is a Scott upset, $\min F\subseteq Y$, and hence $F=\mathord{\uparrow}\min F=\mathord{\uparrow}\operatorname{cl}(F\cap Y)=\mathord{\uparrow}N$. ∎
Remark 3.13.
Our proof that Scott open filters are admissible relies on Item 2.6(3), which requires the Axiom of Choice. An alternative proof using only the Prime Ideal Theorem can be found in [14, Rem. 5.8].
One of the most studied nuclei is the nucleus of double-negation. For a frame $L$ and $a\in L$, recall that the pseudocomplement of $a$ is given by
$a^{*}=\bigvee\{b\in L\mid b\wedge a=0\}.$
The map $a\mapsto a^{**}$ is the double-negation nucleus, and the corresponding sublocale
$\mathfrak{B}(L)\coloneqq\{a\in L\mid a=a^{**}\}$
is the Booleanization of $L$ (see, e.g., [32, p. 246]).
If $j$ is the double-negation nucleus, then $j$-dense elements are simply called dense (see, e.g., [38, p. 131]). It is well known that the corresponding admissible filter dually corresponds to $\max X$, and so we have:
Proposition 3.14.
Let $j$ be the double-negation nucleus on $L$.
  1. (1)
    $H_{j}=\max X$.
  2. (2)
    $N_{j}=\max X$.
Proof.
For (1) see, e.g., [5, Sec. 3]. For (2) observe that (1) and Theorem 3.11 yield $\max X=H_{j}=\mathord{\uparrow}N_{j}$. Consequently, $N_{j}=\max X$. ∎
The following definition is well known. For parts (1) and (2) see, e.g., [25, p. 50] and for part (3) see, e.g., [3, p. 108].
Definition 3.15.
Let $L$ be a frame and $X$ its Priestley space.
  1. (1)
    $j\in N(L)$ is dense if $j0=0$.
  2. (2)
    $S\in S(L)$ is dense if $0\in S$.
  3. (3)
    $N\in N(X)$ is cofinal if $\max X\subseteq N$.
Lemma 3.16.
Let $L$ be a frame, $X$ its Priestley space, and $j\in N(L)$. The following are equivalent.
  1. (1)
    $j$ is dense.
  2. (2)
    $S_{j}$ is dense.
  3. (3)
    $N_{j}$ is cofinal.
Proof.
[…]
Theorem 3.17.
Let $X$ be an L-space. Then $\max X$ is the least cofinal nuclear subset.
Proof.
By Item (2), $\max X$ is a nuclear subset of $X$, and clearly it is the least such containing $\max X$. Thus, it is the least cofinal nuclear subset of $X$. ∎
As a consequence, we obtain the following well-known result of Isbell (see, e.g., [31, p. 40]):
Corollary 3.18 (Isbell’s Density Theorem).
For a frame $L$, the Booleanization $\mathfrak{B}(L)$ is the least dense sublocale of $L$.
Proof.
Let $S_{j}\subseteq L$ be a dense sublocale. Therefore, $N_{j}$ is cofinal by Lemma 3.16, and so $\max X\subseteq N_{j}$ by Theorem 3.17. Consequently, $\mathfrak{B}(L)\subseteq S_{j}$ by Corollaries 3.6 and (2). ∎
4. Priestley duality for the $d$-nucleus
In this section we describe Priestley duals of inductive nuclei, introduced and studied by Martinez and Zenk [29]. We use the Priestley duality tools from previous sections to study the most prominent inductive nucleus, known as the $d$-nucleus. Among other things, we characterize the nuclear set $N_{d}$ corresponding to the $d$-nucleus, and its localic part $Y_{d}$.
Definition 4.1 ([29, Sec. 4]).
A nucleus $j$ on an algebraic frame $L$ is inductive if for all $a\in L$ we have
[…]
Let $X$ be the Priestley space of $L$. As we pointed out after Remark 3.5, we identify $N(L)$ with $N({\sf ClopUp}(X))$, so for each $j\in N(L)$ there is a unique $N\in N(X)$ such that $j\,U=X\setminus\mathord{\downarrow}(N\setminus U)$ for each $U\in{\sf ClopUp}(X)$.
Definition 4.2.
Let $X$ be an L-space. For $j\in N({\sf ClopUp}(X))$ and $U\in{\sf ClopUp}(X)$, define the $j$-core of $U$ by
${\operatorname{core}_{j}}U=\bigcup\{j\,V\mid V\in{\sf ClopSUp}(X)\mbox{ and }V\subseteq U\}.$
Remark 4.3.
Let $j$ be a nucleus on an arithmetic frame $L$ and let $X$ be the Priestley space of $L$. Then $X$ is an arithmetic L-space by Theorem 2.16. Therefore, for all clopen upsets $U,V$ of $X$,
$\operatorname{core}_{j}U\cap\operatorname{core}_{j}V$ $=\bigcup\{j\,U^{\prime}\mid U^{\prime}\in{\sf ClopSUp}(X),\ U^{\prime}\subseteq U\}$
$\qquad\qquad\cap\bigcup\{j\,V^{\prime}\mid V^{\prime}\in{\sf ClopSUp}(X),\ V^{\prime}\subseteq U\}$
$=\bigcup\{j\,(U^{\prime}\cap V^{\prime})\mid U^{\prime},V^{\prime}\in{\sf ClopSUp}(X),\ U^{\prime}\subseteq U,\,V^{\prime}\subseteq V\}$
$=\bigcup\{j\,(W)\mid W\in{\sf ClopSUp}(X),\ W\subseteq U\cap V\}$
$=\operatorname{core}_{j}(U\cap V),$
where the second to last equality is a consequence of $X$ being an arithmetic L-space (see Item 2.15(1)).
Definition 4.4.
Let $X$ be an L-space. We call $N\in N(X)$ inductive if $\mathord{\uparrow}(F\cap N)$ is a Scott upset for each Scott upset $F$ of $X$.
As the name suggests, a nuclear subset is inductive iff its corresponding nucleus is inductive. To prove this, we recall the following:
Lemma 4.5.
Let $X$ be an L-space and $Y$ its localic part.
  1. (1)
    ([11, Lem. 5.1])A closed upset $F$ of $X$ is a Scott upset iff $F\subseteq\operatorname{cl}U$ implies $F\subseteq U$ for all open upsets $U$ of $X$.
  2. (2)
    ([12, Lem. 4.14])Let $y\in Y$ and $U$ an open upset of $X$. Then $y\in U$ iff $y\in\operatorname{cl}U$.
Theorem 4.6.
Let $X$ be an algebraic L-space and $N\in N(X)$. The following are equivalent.
  1. (1)
    NN is inductive.
  2. (2)
    $j_{N}\,U=\operatorname{cl}\operatorname{core}_{j_{N}}U$, for all $U\in{\sf ClopUp}(X)$.
  3. (3)
    $j_{N}$ is inductive.
Proof.
[…]
$\mathord{\uparrow}(\mathord{\uparrow}y\cap N)\subseteq U=\operatorname{cl}\operatorname{core}U,$
and hence $\mathord{\uparrow}(\mathord{\uparrow}y\cap N)\subseteq\operatorname{core}U$ by Item (1). Therefore, since finite unions of clopen Scott upsets are clopen Scott upsets, by compactness there exists $V\in{\sf ClopSUp}(X)$ such that $\mathord{\uparrow}y\cap N\subseteq V$ and $V\subseteq U$. Thus, $y\in j_{N}\,V$ by Item (2), and so $y\in\operatorname{core}_{j_{N}}U$, proving that $j_{N}(U)\cap Y\subseteq\operatorname{core}_{j_{N}}U$.
[…]
$\mathord{\uparrow}y\cap N\subseteq\mathord{\uparrow}(F\cap N)\subseteq U^{\prime},$
so $y\in j_{N}\,U^{\prime}$ by Item (2). Thus, $y\in\operatorname{cl}\operatorname{core}_{j_{N}}U^{\prime}$ by (2), and hence $y\in\operatorname{core}_{j_{N}}U^{\prime}$} by Item (2). Therefore, there exists $V_{y}\in{\sf ClopSUp}(X)$ such that $y\in j_{N}\,V_{y}$ and $V_{y}\subseteq U^{\prime}$}. We have $F=\mathord{\uparrow}\min F\subseteq\bigcup\{j_{N}\,V_{y}\mid y\in\min F\}$, so by compactness there are $V_{1},\dots,V_{n}\in{\sf ClopSUp}(X)$ such that $F\subseteq j_{N}\,V_{1}\cup\dots\cup j_{N}\,V_{n}$. Let $V=V_{1}\cup\dots\cup V_{n}$. Then $V\in{\sf ClopSUp}(X)$. Furthermore, $F\subseteq j_{N}V$ since $j_{N}$ is order-preserving, and clearly $V\subseteq U^{\prime}$. The latter together with Item (1) yields that $V\subseteq U$. Since $F\subseteq j_{N}\,V$, we have $F\cap N\subseteq j_{N}(V)\cap N=V\cap N$ by Item (1). Consequently, $\mathord{\uparrow}(F\cap N)\subseteq U$, and hence $\mathord{\uparrow}(F\cap N)$ is a Scott upset by Item (1).
[…]
$j_{N}\mbox{ is inductive}\iff j_{N}\,U=\bigvee\{j_{N}\,V\mid V\in K({\sf ClopUp}(X))\mbox{ and }V\leq U\}.$
But $K({\sf ClopUp}(X))={\sf ClopSUp}(X)$ (see Remark 2.14) and $\bigvee\mathcal{U}=\operatorname{cl}\bigcup\mathcal{U}$ for $\mathcal{U}\subseteq{\sf ClopUp}(X)$ (see Item 2.6(6)). Therefore,
$\bigvee\{j_{N}\,V\mid V\in K({\sf ClopUp}(X))\mbox{ and }V\leq U\}=\operatorname{cl}\operatorname{core}_{j_{N}}U.$
Consequently, $j_{N}$ is inductive iff $U=\operatorname{cl}\operatorname{core}_{j_{N}}U$ for all $U\in{\sf ClopUp}(X)$. ∎
A prominent example of an inductive nucleus is the so-called $d$-nucleus, introduced by Martinez and Zenk [29, Sec. 5] as a frame-theoretic tool to study $d$-ideals of Riesz spaces (see [22] and [29, Rem. 5.6]).
Definition 4.7.
Let $L$ be an algebraic frame.
  1. (1)
    Define $d:L\to L$ by $da=\bigvee\{k^{**}\mid k\in K(L)\mbox{ and }k\leq a\}$ for all $a\in L$.
  2. (2)
    We call $a\in L$ a $d$-element if $da=a$.
We write $L_{d}$ for the fixpoints of $d$.
Theorem 4.8 ([29, Sec. 5]).
Let $L$ be an algebraic frame.
  1. (1)
    $d$ is a closure operator on $L$.
If in addition $L$ is arithmetic, then
  1. (2)
    $d$ is an inductive dense nucleus on $L$.
  2. (3)
    $L_{d}$ is a sublocale of $L$ and is an arithmetic frame.
We next describe the nuclear subset $N_{d}$ of the Priestley space of $L$ corresponding to the sublocale $L_{d}$. As in Item (1), for each nucleus $j$ on a frame $L$, we view the corresponding nuclear subset $N_{j}$ as the Priestley space of the sublocale $L_{j}$.
Proposition 4.9.
Let $L$ be a frame, $X$ its Priestley space, and $Y$ the localic part of $X$.
  1. (1)
    If $j\in N(L)$, then $N_{j}\cap Y$ is the localic part of $N_{j}$.
If in addition $L$ is an arithmetic frame, then
  1. (2)
    $N_{d}$ is a cofinal inductive nuclear subset of $X$.
  2. (3)
    $N_{d}$ is an arithmetic L-space.
  3. (4)
    $N_{d}=\operatorname{cl}(N_{d}\cap Y)$.
Proof.
(1) Let $Y_{j}$ be the localic part of $N_{j}$. We need to show that $Y_{j}=N_{j}\cap Y$. It is straightforward to see that $N_{j}\cap Y\subseteq Y_{j}$. It remains to show that $Y_{j}\subseteq Y$. Suppose $y\in Y_{j}$. Then $\mathord{\downarrow}y\cap N_{j}$ is clopen in $N_{j}$. Therefore, there is clopen $U\subseteq X$ such that $\mathord{\downarrow}y\cap N_{j}=U\cap N_{j}$. Since $N_{j}$ is a nuclear set, $\mathord{\downarrow}(U\cap N_{j})$ is clopen in $X$. But $\mathord{\downarrow}(U\cap N_{j})=\mathord{\downarrow}(\mathord{\downarrow}y\cap N_{j})=\mathord{\downarrow}y$ because $y\in N_{j}$. Thus, $y\in Y$.
(2) Apply Items (2), 4.6 and 3.16.
(3) Since $L$ is arithmetic, so is $L_{d}$ by Item (3). Therefore, the result follows from Theorem 2.16 because $N_{d}$ is the Priestley dual of $L_{d}$.
(4) $N_{d}\cap Y$ is the localic part of $N_{d}$ by (1), and $N_{d}$ is an SL-space by (3). Hence, $\operatorname{cl}(N_{d}\cap Y)=N_{d}$. ∎
Let $X$ be an arithmetic L-space. Since ${\sf ClopSUp}(X)=K({\sf ClopUp}(X))$ (see Remark 2.14), $d:{\sf ClopUp}(X)\to{\sf ClopUp}(X)$ is given by
$d\,U=\operatorname{cl}\bigcup\{V^{**}\mid V\in{\sf ClopSUp}(X)\mbox{ and }V\subseteq U\},$
where $V^{*}=X\setminus\mathord{\downarrow}V$ (see, e.g., [20, p. 20]), so $x\in V^{**}$ iff $\mathord{\uparrow}x\subseteq\mathord{\downarrow}V$. We also recall (see Definition 4.2) that the $d$-core of $U\in{\sf ClopUp}(X)$ is given by
$\operatorname{core}_{d}U=\bigcup\{d\,V\mid V\in{\sf ClopSUp}(X),\,V\subseteq U\}.$
Lemma 4.10.
Let $L$ be an arithmetic frame, $X$ its Priestley space, and $Y$ the localic part of $X$.
  1. (1)
    If $U\in{\sf ClopSUp}(X)$, then $d\,U=U^{**}$.
  2. (2)
    If $U\in{\sf ClopUp}(X)$, then $x\in\operatorname{core}_{d}U$ iff $\mathord{\uparrow}x\subseteq\mathord{\downarrow}\operatorname{core}U$.
Proof.
(1) $d\,U=\operatorname{cl}\bigcup\{V^{**}\mid V\in{\sf ClopSUp}(X)\mbox{ and }V\subseteq U\}=U^{**}$ since $U\in{\sf ClopSUp}(X)$.
(2) First suppose that $x\in\operatorname{core}_{d}U$. Then there is $V\in{\sf ClopSUp}(X)$ with $x\in d\,V$ and $V\subseteq U$. Therefore, $x\in V^{**}$ by (1), which means that $\mathord{\uparrow}x\subseteq\mathord{\downarrow}V$. Since $V$ is a Scott upset, $V\subseteq U$ implies $V\subseteq\operatorname{core}U$. Thus, $x\in d\,V$ implies $\mathord{\uparrow}x\subseteq\mathord{\downarrow}\operatorname{core}U$. For the converse, if $\mathord{\uparrow}x\subseteq\mathord{\downarrow}\operatorname{core}U$, then
$\mathord{\uparrow}x\subseteq{\big\downarrow}\bigcup\{V\in{\sf ClopSUp}(X)\mid V\subseteq U\}=\bigcup\{\mathord{\downarrow}V\mid V\subseteq U\mbox{ and }V\in{\sf ClopSUp}(X)\}.$
Hence, by Item 2.6(5), $\{\mathord{\downarrow}V\mid V\subseteq U\mbox{ and }V\in{\sf ClopSUp}(X)\}$ is an open cover of $\mathord{\uparrow}x$. Since $\mathord{\uparrow}x$ is compact and this open cover is directed, there is $V\in{\sf ClopSUp}(X)$ such that $V\subseteq U$ and $\mathord{\uparrow}x\subseteq\mathord{\downarrow}V$. This yields that $x\in V^{**}$, so $x\in d\,V$ by (1). Consequently, $x\in\operatorname{core}_{d}U$. ∎
Let $X$ be an arithmetic L-space and $Y$ its localic part. We let $Y_{d}$ denote the localic part of $N_{d}$. By Item (1), $Y_{d}=N_{d}\cap Y$. We conclude the section by giving several characterizations of $Y_{d}$. For this we need the following lemma.
Lemma 4.11.
Let $X$ be an algebraic L-space and $F$ a Scott upset of $X$. Then
$F=\bigcap\{U\in{\sf ClopSUp}(X)\mid F\subseteq U\}.$
Proof.
Since $F$ is a closed upset, $F=\bigcap\{U\in{\sf ClopUp}(X)\mid F\subseteq U\}$ (see Item 2.6(2)). Thus, it suffices to show that for each $U\in{\sf ClopUp}(X)$ with $F\subseteq U$, there is $V\in{\sf ClopSUp}(X)$} with $F\subseteq V\subseteq U$. Since $X$ is an algebraic L-space, $U=\operatorname{cl}\operatorname{core}U$, so $F\subseteq U$ implies that $F\subseteq\operatorname{core}U$ by Item (1). Now apply compactness to obtain the desired $V$. ∎
Theorem 4.12.
Let $L$ be an arithmetic frame, $X$ its Priestley space, and $Y$ the localic part of $X$. For $y\in Y$, the following are equivalent.
  1. (1)
    $y\in Y_{d}$.
  2. (2)
    $\forall U\in{\sf ClopUp}(X),\,y\in\operatorname{core}_{d}U\implies y\in U$.
  3. (3)
    $\forall V\in{\sf ClopSUp}(X),\,\max\mathord{\uparrow}y\subseteq V\implies y\in V$.
  4. (4)
    $\{y\}=\max(\mathord{\downarrow}x\cap Y)$ for some $x\in\max X$.
Proof.
[…]
5. $\max Y$ and regularity of $L_{d}$
Martinez and Zenk [29, Prop. 5.2] characterized when $L_{d}$ is a regular frame. In this section we give several alternative characterizations, utilizing Priestley duality. This, in particular, involves the maximal spectrum $\max Y$ of the localic part $Y$ of the Priestley space of $L$. As a consequence, we obtain that $L_{d}$ is regular iff $L_{d}$ is locally Stone.
Recall (see, e.g., [31, p. 89]) that a frame $L$ is regular if for all $a\in L$ we have
$a=\bigvee\{b\in L\mid b^{*}\vee a=1\}.$
Priestley spaces of regular frames were studied in [36, 7, 12]. We recall:
Definition 5.1 ([12, Defs. 7.1 and 7.6]).
Let $X$ be an L-space.
  1. (1)
    For $U\in{\sf ClopUp}(X)$, the regular part of $U$ is
    $\operatorname{reg}U=\bigcup\{V\in{\sf ClopUp}(X)\mid\mathord{\downarrow}V\subseteq U\}.$
  2. (2)
    $X$ is L-regular if $\operatorname{cl}\operatorname{reg}U=U$ for each $U\in{\sf ClopUp}(X)$.
Theorem 5.2.
Let $L$ be a frame, $X$ its Priestley space, and $Y$ the localic part of $X$.
  1. (1)
    [7, Lem 3.6] $L$ is regular iff $X$ is L-regular.
  2. (2)
    [12, Proof of Thm. 7.11] If $X$ is an SL-space and $Y$ is regular, then $X$ is L-regular.
  3. (3)
    [12, Lem. 7.15(3)] If $X$ is L-regular, then $Y\subseteq\min X$.
An element $p\neq 1$ of a frame $L$ is (meet-)prime if $a\wedge b\leq p$ implies $a\leq p$ or $b\leq p$ (see, e.g., [31, p. 13]). A prime element $p$ is minimal prime with respect to $a\in L$ if $p$ is minimal among the primes $q\geq a$. It is known (see, e.g., [30, p. 264]) that every prime element $p$ greater than $a\in L$ has a minimal prime element $q$ with respect to $a$ beneath it. Since the assignment $p\mapsto L\setminus\mathord{\downarrow}p$ establishes an isomorphism between the posets of prime elements and completely prime filters (see, e.g., [31, p. 14]), this condition can equivalently be formulated as follows: for every completely prime filter $P$ contained in a filter $F$, there exists a completely prime filter $Q$ that is maximal among the completely prime filters contained in $F$. Thus, we arrive at the following lemma, which gives the means to find (relatively) maximal localic points.
Lemma 5.3.
Let $L$ be a frame, $X$ its Priestley space, $Y$ the localic part of $X$, and $y\in Y$.
  1. (1)
    $\mathord{\uparrow}y\cap\max Y\neq\varnothing$.
  2. (2)
    $\mathord{\uparrow}y\cap\max(\mathord{\downarrow}x\cap Y)\neq\varnothing$ for every $x\in X$ with $y\leq x$.
Proof.
For (1) take $P=y$ and $F=L$, and for (2) take $P=y$ and $F=x$. ∎
We show that $\max Y\subseteq Y_{d}$, but that the converse is not true in general. For this we require the following lemma.
Lemma 5.4.
Let $L$ be an arithmetic frame, $X$ its Priestley space, and $Y$ the localic part of $X$.
  1. (1)
    If $F$ and $G$ are Scott upsets of $X$, then so is $F\cap G$.
  2. (2)
    Suppose $y\in\max Y$ and $F$ is a Scott upset of $X$. If $\mathord{\uparrow}y\cap F\neq\varnothing$, then $y\in F$.
Proof.
(1) This can be seen by applying [13, Lem. 5.2] and [12, Lem. 6.3(2)]. To keep the proof self-contained, we give a short argument. By Lemma 4.11, $F$ and $G$ are intersections of down-directed families of clopen Scott upsets. Therefore, so is $F\cap G$ since the binary intersection of clopen Scott upsets is a clopen Scott upset ($X$ is an arithmetic L-space). Thus, $F\cap G$ is a Scott upset because the intersection of a down-directed family of clopen Scott upsets is a Scott upset (see [12, Lem. 5.14(1)]).
(2) Since $\mathord{\uparrow}y\cap F$ is a nonempty closed upset, $\min(\mathord{\uparrow}y\cap F)\neq\varnothing$ (see, e.g., [20, Thm. 3.2.1]). Because $\mathord{\uparrow}y,F$ are Scott upsets, $\mathord{\uparrow}y\cap F$ is also a Scott upset by (1). Therefore, $\min(\mathord{\uparrow}y\cap F)\subseteq Y$. Since $y$ is underneath each point in $\min(\mathord{\uparrow}y\cap F)$ and $y\in\max Y$, we conclude that $\min(\mathord{\uparrow}y\cap F)=\{y\}$, and hence $y\in F$. ∎
Theorem 5.5.
Let $X$ be an algebraic L-space and $Y$ its localic part.
  1. (1)
    If $N\subseteq X$ is a cofinal inductive nuclear subset, then $\max Y\subseteq N$.
  2. (2)
    If $X$ is the Priestley space of an arithmetic frame $L$, then $\max Y\subseteq Y_{d}$.
Proof.
(1) Let $y\in\max Y$. Since $\mathord{\uparrow}y$ is a Scott upset and $N$ is inductive, $\mathord{\uparrow}(\mathord{\uparrow}y\cap N)$ is a Scott upset. Because $N$ is cofinal, $\max X\subseteq N$, and thus $\mathord{\uparrow}y\cap N\neq\varnothing$. Therefore, $\mathord{\uparrow}(\mathord{\uparrow}y\cap N)\subseteq\mathord{\uparrow}y$ is a nonempty Scott upset, so $y\in\mathord{\uparrow}(\mathord{\uparrow}y\cap N)$ by Item (2). Hence, $\mathord{\uparrow}y\subseteq\mathord{\uparrow}(\mathord{\uparrow}y\cap N)$. Consequently, $\mathord{\uparrow}y=\mathord{\uparrow}(\mathord{\uparrow}y\cap N)$, and so $y\in N$.
(2) By Item (2), $N_{d}$ is a cofinal inductive nuclear subset of $X$. Therefore, $\max Y\subseteq N_{d}$ by (1), and so $\max Y\subseteq Y_{d}$ by Item (1). ∎
Example 5.6.
To see that in general $\max Y\neq Y_{d}$, let $\beta\mathbb{N}$ be the Stone-Čech compactification of the natural numbers (see, e.g., [18, p. 174]). As is customary, we write $\mathbb{N}^{*}$ for the remainder. Let $X=\beta\mathbb{N}\cup\{y\}$, where the order on $X$ is defined as shown in Figure 1(a). It is well known (see, e.g., [17, p. 28]) that $\beta\mathbb{N}$ is homeomorphic to the Stone space of the powerset $\wp(\mathbb{N})$ of $\mathbb{N}$. Therefore, $X$ is homeomorphic to the Priestley space of the lattice $L$ obtained by adding a new top to $\wp(\mathbb{N})$; see Figure 1(b). Since $\wp(\mathbb{N})$ is an arithmetic frame and $1\in K(L)$, it is clear that so is $L$, and hence $X$ is an arithmetic L-space.
Because the set of isolated points of $X$ is $\mathbb{N}\cup\{y\}$, we have that $\mathord{\downarrow}x$ is clopen iff $x$ is an isolated point of $X$. Thus, the localic part of $X$ is $Y=\mathbb{N}\cup\{y\}$. Therefore, $\max Y=\mathbb{N}$. On the other hand, $y\in N_{d}$ by Item (4), so $Y_{d}=Y$. Consequently, $Y_{d}\neq\max Y$.
[…]
(a) The Priestley space $\beta\mathbb{N}\cup\{y\}.$
$\wp(\mathbb{N})$
(b) The lattice $\wp(N)$ with a new top.
Figure 1. An arithmetic L-space in which $Y_{d}\neq\max Y$.
In order to characterize when $L_{d}$ is regular, we require the following lemma, which allows us to separate Scott upsets from maximal localic points via clopen Scott upsets.
Lemma 5.7.
Let $X$ be an arithmetic L-space, $Y$ its localic part, and $F$ a Scott upset of $X$. If $y\in\max Y\setminus F$, then there is $U\in{\sf ClopSUp}(X)$ with $y\in U$ and $U\cap F=\varnothing$.
Proof.
Since $y\not\in F$, we have $\mathord{\uparrow}y\cap F=\varnothing$ by Item (2). Therefore, Lemma 4.11 yields that $\bigcap\{U\in{\sf ClopSUp}(X)\mid y\in U\}\cap F=\varnothing$. Thus, we can use compactness and the fact that finite intersections of clopen Scott upsets are again Scott upsets ($X$ is an arithmetic L-space) to produce $U\in{\sf ClopSUp}(X)$ with $y\in U$ and $U\cap F=\varnothing$. ∎
We recall that a space $X$ is locally Stone if $X$ is zero-dimensional, locally compact, and Hausdorff. Thus, $X$ is locally Stone if in the definition of a Stone space we weaken compactness to local compactness. A frame $L$ is locally Stone if it is isomorphic to the frame of opens of a locally Stone space. By [10, Thm. 3.11], a frame is locally Stone iff it is algebraic and zero-dimensional (each element is a join of complemented elements). We will use the following fact: if $L$ is an algebraic frame and $X$ its Priestley space, then the localic part $Y$ of $X$ is locally compact (see [12, Thm. 5.10], where the result is proved in the more general setting of continuous frames).
Theorem 5.8.
Let $L$ be an arithmetic frame, $X$ its Priestley space, and $Y$ the localic part of $X$. The following are equivalent.
  1. (1)
    $L_{d}$ is regular.
  2. (2)
    $N_{d}$ is L-regular.
  3. (3)
    $Y_{d}$ is an antichain.
  4. (4)
    $\max Y=Y_{d}$.
  5. (5)
    $Y_{d}$ is a locally Stone space.
  6. (6)
    $L_{d}$ is a locally Stone frame.
Proof.
[…]
$U\cap\max Y=\operatorname{cl}(\operatorname{core}U)\cap\max Y=\operatorname{core}U\cap\max Y.$
Thus, to see that $Y_{d}$ is zero-dimensional, it is enough to show that $U\cap\max Y$ is clopen for each $U\in{\sf ClopSUp}(X)$. For this it is sufficient to show that for each $y\in\max Y\setminus U$ there is $V\in{\sf ClopSUp}(X)$ with $y\in V\cap\max Y\subseteq\max Y\setminus U$. But this follows from Lemma 5.7. Finally, to see that $Y_{d}$ is Hausdorff, let $y,y^{\prime}\in Y_{d}=\max Y$ be distinct. Then $y\not\in\mathord{\uparrow}y^{\prime}$, so by Lemma 5.7 there is $U\in{\sf ClopSUp}(X)$ such that $y\in U$ and $y^{\prime}\not\in U$. But then $y\in U\cap Y_{d}$, which is clopen by the above.
[…]
6. Spectra of maximal $d$-elements
In this section, we begin our investigation of the spectrum $\max L_{d}$ of maximal $d$-elements of an arithmetic frame $L$, as introduced in [8]. First, we show that $\max L_{d}$ is in a bijective correspondence with $\min Y_{d}$. Following this, we establish that $\min Y_{d}$, viewed as a subspace of $Y$, is homeomorphic to $\max L_{d}$. The homeomorphism enables us to analyze the properties of $\max L_{d}$ through $\min Y_{d}$. We show that the frame $\Omega(\min Y_{d})$ of open subsets of $\min Y_{d}$ can be realized as a sublocale of $L$, and describe the corresponding nuclear subset of $X$. We conclude the section by observing that the localic part of this nuclear subset is the soberification of $\min Y_{d}$.
Definition 6.1.
Let $X$ be an arithmetic L-space and $U\in{\sf ClopUp}(X)$.
  1. (1)
    We call $U$ a $d$-upset if $\operatorname{cl}\operatorname{core}_{d}U=U$.
  2. (2)
    We call $U$ a maximal $d$-upset if it is maximal among proper $d$-upsets of $X$.
Remark 6.2.
Since $d$ is inductive, it is straightforward to verify that (maximal) $d$-elements of an arithmetic frame correspond to (maximal) $d$-upsets of its Priestley space.
We now show that maximal $d$-upsets are in one-to-one correspondence with elements of $\min Y_{d}$. For this, we require the following lemmas.
Lemma 6.3.
Let $X$ be an arithmetic L-space and $U\in{\sf ClopUp}(X)$. Then $\operatorname{cl}\operatorname{core}_{d}U=X$ iff $Y_{d}\subseteq U$.
Proof.
Since $d$ is inductive, $\operatorname{cl}\operatorname{core}_{d}U=X$ iff dU=Xd\,U=X (see Item (2)). It follows from Item (5) that dU=Xd\,U=X iff $N_{d}\subseteq U$. But by Item (4), $N_{d}=\operatorname{cl}Y_{d}$, so $N_{d}\subseteq U$ iff $Y_{d}\subseteq U$ since $U$ is closed. ∎
Lemma 6.4.
Let $X$ be an arithmetic L-space and $y\in Y_{d}$.
  1. (1)
    $X\setminus\mathord{\downarrow}y$ is a $d$-upset.
  2. (2)
    $X\setminus\mathord{\downarrow}y$ is a maximal $d$-upset iff $y\in\min Y_{d}$.
  3. (3)
    Maximal $d$-upsets are precisely the clopen upsets of the form $X\setminus\mathord{\downarrow}y$ for some $y\in\min Y_{d}$}.
Proof.
(1) We need to show $X\setminus\mathord{\downarrow}y=\operatorname{cl}\operatorname{core}_{d}(X\setminus\mathord{\downarrow}y)$. For this it is sufficient to show that $Y\cap(X\setminus\mathord{\downarrow}y)=Y\cap\operatorname{core}_{d}(X\setminus\mathord{\downarrow}y)$} since $X$ is a spatial L-space. We have the inclusion $Y\cap(X\setminus\mathord{\downarrow}y)\subseteq Y\cap\operatorname{core}_{d}(X\setminus\mathord{\downarrow}y)$} since $Y\cap U=Y\cap\operatorname{core}U$ and $\operatorname{core}U\subseteq\operatorname{core}_{d}U$ for each clopen upset $U$. For the reverse inclusion, let $z\in Y$ and suppose towards a contradiction that $z\in\operatorname{core}_{d}(X\setminus\mathord{\downarrow}y)$ and $z\not\in X\setminus\mathord{\downarrow}y$. Then $z\leq y$, so $y\in\operatorname{core}_{d}(X\setminus\mathord{\downarrow}y)$. Therefore, $y\in X\setminus\mathord{\downarrow}y$ by Theorem 4.12 ($y\in Y_{d}$), a contradiction.
(2) Suppose $y\not\in\min Y_{d}$. Then there exists $y^{\prime}\in Y_{d}$ with $y^{\prime}<y$. Therefore, there exists $V\in{\sf ClopUp}(X)$ containing $y$ and missing $y^{\prime}$. Let $U=\operatorname{cl}\operatorname{core}_{d}(V\cup(X\setminus\mathord{\downarrow}y))$. Then $U$ is a $d$-upset and $X\setminus\mathord{\downarrow}y\subsetneq U$ because $y\in U$ and $y\notin X\setminus\mathord{\downarrow}y$. But $U=\operatorname{cl}\operatorname{core}_{d}(V\cup(X\setminus\mathord{\downarrow}y))\neq X$ by Lemma 6.3 since $y^{\prime}\in Y_{d}$ and $y^{\prime}\notin V\cup(X\setminus\mathord{\downarrow}y)$, yielding that $Y_{d}\not\subseteq V\cup(X\setminus\mathord{\downarrow}y)$. Thus, $X\setminus\mathord{\downarrow}y$ is not a maximal $d$-upset.
Suppose $y\in\min Y_{d}$ and $X\setminus\mathord{\downarrow}y\subsetneq U$ for a $d$-upset $U$. Then $y\in U$, and since $y\in\min Y_{d}$ we get $Y_{d}\subseteq(X\setminus\mathord{\downarrow}y)\cup\{y\}\subseteq U$. Therefore, $U=\operatorname{cl}\operatorname{core}_{d}U=X$ by Lemma 6.3.
(3) By (2) it suffices to show that every maximal $d$-upset is of the desired form, so suppose $U\in{\sf ClopUp}(X)$ is a maximal $d$-upset. Then $\operatorname{cl}\operatorname{core}_{d}U=U\neq X$, so $Y_{d}\not\subseteq U$ by Lemma 6.3. Therefore, there exists $y\in Y_{d}\setminus U$. Then $\mathord{\downarrow}y\cap U=\varnothing$, so $U\subseteq X\setminus\mathord{\downarrow}y\neq X$. But $X\setminus\mathord{\downarrow}y$ is a $d$-upset by (1). Hence, $U=X\setminus\mathord{\downarrow}y$ since $U$ is a maximal $d$-upset. Thus, $y\in\min Y_{d}$ by (2). ∎
As an immediate consequence, we obtain:
Theorem 6.5.
Let $X$ be an arithmetic L-space. The map $y\mapsto X\setminus\mathord{\downarrow}y$ is a bijection from $\min Y_{d}$ to the collection of maximal $d$-upsets of $X$.
Equipping $Y_{d}$ with the subspace topology inherited from $Y$, we have:
Theorem 6.6.
$\min Y_{d}$ is homeomorphic to $\max L_{d}$.
Proof.
Since $\varphi:L\to{\sf ClopUp}(X)$ is an isomorphism, define $\alpha:\min Y_{d}\to\max L_{d}$ by $\alpha(y)=\varphi^{-1}(X\setminus\mathord{\downarrow}y)$. By Theorem 6.5, $\alpha$ is a bijection. Thus, it suffices to show that for all $U\subseteq\min Y_{d}$ we have $U$ is open iff $\alpha(U)$ is open. Now, $U$ is open iff $U=V\cap\min Y_{d}$ for some $V\in{\sf ClopUp}(X)$, and $\alpha(U)$ is open iff $\alpha(U)=\{m\in\max L_{d}\mid a\not\leq m\}$ for some $a\in L$ (see, e.g., [8, Sec. 3]). Since $m\in\max(L_{d})$ iff $\varphi(m)$ is a maximal $d$-upset, by Item (3) we have that $m\in\max(L_{d})$ iff $\varphi(m)=X\setminus\mathord{\downarrow}y$ for some $y\in\min Y_{d}$. Moreover, for $a\in L$, we have that $\varphi(a)\not\subseteq X\setminus\mathord{\downarrow}y$ iff $y\in\varphi(a)$. Therefore,
$\alpha(U)=\{m\in\max L_{d}\mid a\not\leq m\}$ $\iff\varphi[\alpha(U)]=\{\varphi(m)\mid\varphi(a)\not\subseteq\varphi(m)\}$
$\iff\varphi[\alpha(U)]=\{X\setminus\mathord{\downarrow}y\mid y\in\min Y_{d},\ \varphi(a)\not\subseteq X\setminus\mathord{\downarrow}y\}$
$\iff U=\{y\in\min Y_{d}\mid y\in\varphi(a)\}$
$\iff U=\varphi(a)\cap\min Y_{d}.\qed$
As we pointed out in the introduction, if $L$ is an arithmetic frame with a unit (see the beginning of Section 7), then $\max L_{d}$ is a compact $T_{1}$-space. We next show that being $T_{1}$ does not depend on the existence of a unit.
Proposition 6.7.
$\min Y_{d}$ is $T_{1}$.
Proof.
Suppose $y,y^{\prime}\in\min Y_{d}$ are distinct. Then $y\not\leq y^{\prime}$ since $\min Y_{d}$ is an antichain. By Priestley separation, there exists $U\in{\sf ClopUp}(X)$ with $y\in U$ and $y^{\prime}\notin U$. Hence, $U\cap\min Y_{d}$ is an open subset of $\min Y_{d}$ containing $y$ and missing $y^{\prime}$. ∎
We now concentrate on the frame $\Omega(\min Y_{d})$ of open subsets of $\min Y_{d}$ and show that it can be realized as a sublocale of $L$. To this end, since $L$ is isomorphic to ${\sf ClopUp}(X)$, we introduce a nucleus on ${\sf ClopUp}(X)$ that determines $\Omega(\min Y_{d})$.
Lemma 6.8.
The map $h:{\sf ClopUp}(X)\to\Omega(\min Y_{d})$, given by $h(U)=U\cap\min Y_{d}$, is an onto frame homomorphism.
Proof.
It is clear that $h$ is onto and preserves finite meets. To see that it preserves arbitrary joins let $\{U_{i}\}\subseteq{\sf ClopUp}(X)$. Then, by Item (2),
$h\left(\bigvee U_{i}\right)=\left(\operatorname{cl}\bigcup U_{i}\right)\cap\min Y_{d}=\left(\bigcup U_{i}\right)\cap\min Y_{d}=\bigcup(U_{i}\cap\min Y_{d})=\bigcup h(U_{i}).\qed$
By the previous lemma, there is a nucleus $\mathord{\rho}=h_{*}\circ h:{\sf ClopUp}(X)\to{\sf ClopUp}(X)$, where $h_{*}$ is the right adjoint of $h$ (see, e.g., [31, p. 31]). Then, for each $U\in{\sf ClopUp}(X)$,
$\mathord{\rho}(U)$ $=\bigvee\{V\in{\sf ClopUp}(X)\mid h(V)\subseteq h(U)\}$
$=\operatorname{cl}\bigcup\{V\in{\sf ClopUp}(X)\mid V\cap\min Y_{d}\subseteq U\cap\min Y_{d}\}$
$=\operatorname{cl}\bigcup\{V\in{\sf ClopUp}(X)\mid V\cap\min Y_{d}\subseteq U\}.$
Lemma 6.9.
Let $L$ be an arithmetic frame and $X$ its L-space. For $a\in L$ set
$M_{a}=\bigwedge\{m\in\max L_{d}\mid a\leq m\}.$
Then $\varphi\left(\bigwedge M_{a}\right)=\mathord{\rho}(\varphi(a))$.
Proof.
Observe that
$\varphi\left(\bigwedge M_{a}\right)$ $=\varphi\left(\bigvee\{b\in L\mid b\leq m\mbox{ for all }m\in M_{a}\}\right)$
$=\operatorname{cl}\bigcup\left\{V\in{\sf ClopUp}(X)\mid V\subseteq\bigcap\varphi[M_{a}]\right\}.$
Recall that $m\in\max L_{d}$ iff $\varphi(m)$ is a maximal $d$-upset. Thus, using Item (3), we obtain that $m\in M_{a}$ iff $\varphi(m)=X\setminus\mathord{\downarrow}y$ for some $y\in\min Y_{d}\setminus\varphi(a)$. Therefore,
$\bigcap\varphi[M_{a}]=\bigcap\{X\setminus\mathord{\downarrow}y\mid y\in\min Y_{d}\setminus\varphi(a)\}=X\setminus\mathord{\downarrow}(\min Y_{d}\setminus\varphi(a)).$
Consequently, since $V$ is an upset,
$V\subseteq\bigcap\varphi[M_{a}]$ $\iff V\subseteq X\setminus\mathord{\downarrow}(\min Y_{d}\setminus\varphi(a))$
$\iff V\cap\mathord{\downarrow}(\min Y_{d}\setminus\varphi(a))=\varnothing$
$\iff V\cap(\min Y_{d}\setminus\varphi(a))=\varnothing$
$\iff V\cap\min Y_{d}\subseteq\varphi(a),$
and the result follows from the above description of $\rho$. ∎
Remark 6.10.
By the previous lemma, the nucleus $\rho$ can be defined on an arbitrary arithmetic frame $L$ by
$\mathord{\rho}(a)=\bigwedge\{m\in\max L_{d}\mid a\leq m\}$
for each $a\in L$.
We now describe the nuclear subset of $X$ corresponding to the nucleus $\mathord{\rho}$. For this we use the following:
Theorem 6.11.
Let $L$ be an arithmetic $L$-space. Then $N_{\mathord{\rho}}=\operatorname{cl}\min Y_{d}$.
Proof.
By Item (4), $\operatorname{cl}\min Y_{d}$ is a nuclear subset of $X$. Let $j\in N(L)$ be the nucleus associated with $\operatorname{cl}\min Y_{d}$ (see Remark 3.5). It suffices to show that $\varphi(j(a))=\mathord{\rho}(\varphi(a))$ for all $a\in L$.
($\subseteq$) Let $x\in\varphi(j(a))$. Then $\mathord{\uparrow}x\cap\operatorname{cl}\min Y_{d}\subseteq\varphi(a)$ by Item (2). Since $\mathord{\uparrow}x$ is a closed upset, it is an intersection of clopen upsets (see Item 2.6(2)). Therefore, by compactness, there is $V\in{\sf ClopUp}(X)$ such that $x\in V$ and $V\cap\operatorname{cl}\min Y_{d}\subseteq\varphi(a)$. Thus, $V\cap\min Y_{d}\subseteq\varphi(a)$, and so $x\in\rho(\varphi(a))$.
($\supseteq$) By Item (2),
$\rho(\varphi(a))\cap\min Y_{d}$ $=\operatorname{cl}\bigcup\{V\in{\sf ClopUp}(X)\mid V\cap\min Y_{d}\subseteq\varphi(a)\}\cap\min Y_{d}$
$=\bigcup\{V\in{\sf ClopUp}(X)\mid V\cap\min Y_{d}\subseteq\varphi(a)\}\cap\min Y_{d}\subseteq\varphi(a).$
Therefore, since $\varphi(a)$ is closed,
$\rho(\varphi(a))\cap\operatorname{cl}\min Y_{d}\subseteq\operatorname{cl}(\rho(\varphi(a))\cap\min Y_{d})\subseteq\varphi(a).$
Thus, for each $x\in\rho(\varphi(a))$,
$\mathord{\uparrow}x\cap\operatorname{cl}\min Y_{d}\subseteq\rho(\varphi(a))\cap\operatorname{cl}\min Y_{d}\subseteq\varphi(a).$
Consequently, $x\in\varphi(ja)$ by Item (2). ∎
We conclude this section by describing the link between $\min Y_{d}$ and the localic part of $N_{\mathord{\rho}}$.
Proposition 6.12.
The localic part $Y_{\mathord{\rho}}$ of $N_{\mathord{\rho}}$ is the soberification of $\min Y_{d}$.
Proof.
The localic part of any L-space is the space of points of the associated frame (see Remark 2.9). Thus, $Y_{\mathord{\rho}}$ is the space of points of $\Omega(\min Y_{d})$, which is the soberification of $\min Y_{d}$ (see, e.g., [25, p. 44]). ∎
7. Compactness of the maximal $d$-spectrum
We now turn our attention to studying topological properties of $\min Y_{d}$. As we mentioned in the introduction, in [8] the second author only considered arithmetic frames with a unit; that is, a compact dense element, where we recall (see the paragraph before Item (1)) that an element $a\in L$ is dense if $a^{**}=1$. In this section, we characterize units in the language of Priestley spaces and compare the existence of a unit to compactness of $\min Y_{d}$.
We start by characterizing compact subsets of $\min Y_{d}$ in terms of special Scott upsets of $X$.
Definition 7.1.
Let $X$ be an arithmetic L-space. A subset $Z\subseteq X$ is called $d$-initial if $Z\cap Y\subseteq\mathord{\uparrow}(Z\cap\min Y_{d})$.
Lemma 7.2.
Let $X$ be an arithmetic L-space and $Y$ its localic part.
  1. (1)
    A Scott upset $F\subseteq X$ is $d$-initial iff $F=\mathord{\uparrow}(F\cap\min Y_{d})$.
  2. (2)
    If $F\subseteq X$ is a $d$-initial Scott upset, then $F\cap\min Y_{d}$ is compact.
  3. (3)
    If $K\subseteq\min Y_{d}$ is compact then $\mathord{\uparrow}K$ is a $d$-initial Scott upset.
Proof.
(1) The right-to-left implication is immediate. For the left-to-right implication, let $F$ be a $d$-initial Scott upset. Then $\min F\subseteq F\cap Y\subseteq\mathord{\uparrow}(F\cap\min Y_{d})$, and hence $F=\mathord{\uparrow}\min F=\mathord{\uparrow}(F\cap\min Y_{d})$.
(2) Suppose that $F\cap\min Y_{d}\subseteq\bigcup(U_{i}\cap\min Y_{d})$ for a family $\{U_{i}\}\subseteq{\sf ClopUp}(X)$. Then $F\cap\min Y_{d}\subseteq\bigcup U_{i}$, and so $F\subseteq\bigcup U_{i}$ by (1). Since $F$ is closed, it is compact, and hence $F\subseteq U_{i_{1}}\cup\dots\cup U_{i_{n}}$ for some $i_{1},\dots,i_{n}$. Therefore,
$F\cap\min Y_{d}\subseteq(U_{i_{1}}\cap\min Y_{d})\cup\dots\cup(U_{i_{n}}\cap\min Y_{d}),$
and hence $F\cap\min Y_{d}$ is compact.
(3) Clearly, $\mathord{\uparrow}K$ is $d$-initial. Thus, it suffices to show that $\mathord{\uparrow}K$ is a Scott upset. Since $\min\mathord{\uparrow}K=K\subseteq\min Y_{d}$, it is enough to show that $\mathord{\uparrow}K$ is closed. Let $x\notin\mathord{\uparrow}K$. Then $y\not\leq x$ for all $y\in K$. By Priestley separation, there is $U_{y}\in{\sf ClopUp}(X)$ such that $y\in U_{y}$ and $x\not\in U_{y}$. Therefore, $K\subseteq\bigcup(U_{y}\cap\min Y_{d})$, so by compactness of $K$ there is $U\in{\sf ClopUp}(X)$ such that $\mathord{\uparrow}K\subseteq U$ and $x\notin U$. Thus, $\mathord{\uparrow}K$ is closed. ∎
As an immediate consequence, we have:
Proposition 7.3.
Let $X$ be an arithmetic $L$-space and $K\subseteq\min Y_{d}$. The following are equivalent:
  1. (1)
    $K$ is compact.
  2. (2)
    $\mathord{\uparrow}K$ is a $d$-initial Scott upset.
  3. (3)
    There is a $d$-initial Scott upset $F\subseteq X$ such that $K=F\cap\min Y_{d}$.
Theorem 7.4.
Let $X$ be an arithmetic L-space. There is a poset isomorphism between compact subsets of $\min Y_{d}$ and $d$-initial Scott upsets of $X$ (both ordered by inclusion).
Proof.
Consider the maps $F\mapsto F\cap\min Y_{d}$ and $K\mapsto\mathord{\uparrow}K$, where $F\subseteq X$ is a $d$-initial Scott upset and $K\subseteq\min Y_{d}$ is compact. These are well defined by Lemma 7.2, and are clearly order preserving. It suffices to show that these maps are inverses of each other. But $F=\mathord{\uparrow}(F\cap\min Y_{d})$ by Item (1), and it is easy to see that $K=\mathord{\uparrow}K\cap\min Y_{d}$, completing the proof. ∎
We recall (see, e.g., [31, p. 25]) that a frame is max-bounded if each proper element is below a maximal element.
Proposition 7.5.
Let $L$ be an arithmetic frame and $X$ its Priestley space. Then $L_{d}$ is max-bounded iff $N_{d}$ is $d$-initial.
Proof.
Since $L_{d}$ is max-bounded iff every proper $d$-upset is contained in a maximal $d$-upset (see Remark 6.2), it suffices to show that the latter condition is equivalent to $N_{d}$ being $d$-initial.
($\Rightarrow$) Let $y\in N_{d}\cap Y$. Then $y\in Y_{d}$ by Item (1), so $X\setminus\mathord{\downarrow}y$ is a $d$-upset by Item (1). Hence, there exists a maximal $d$-upset $U$ such that $X\setminus\mathord{\downarrow}y\subseteq U$. By Item (3), $U=X\setminus\mathord{\downarrow}y^{\prime}$ for some $y^{\prime}\in\min Y_{d}$. Thus, $X\setminus\mathord{\downarrow}y\subseteq X\setminus\mathord{\downarrow}y^{\prime}$, which implies that $y^{\prime}\in\mathord{\downarrow}y^{\prime}\subseteq\mathord{\downarrow}y$. Therefore, $y^{\prime}\leq y$, as required.
($\Leftarrow$) Let $U$ be a proper $d$-upset. Then $U=\operatorname{cl}\operatorname{core}_{d}U\neq X$, so $Y_{d}\not\subseteq U$ by Lemma 6.3. Hence, there is $y\in Y_{d}\setminus U\subseteq N_{d}$. Since $N_{d}$ is $d$-initial, there is $y^{\prime}\in\min Y_{d}$ with $y^{\prime}\leq y$. Therefore, $U\subseteq X\setminus\mathord{\downarrow}y^{\prime}$, which is a maximal $d$-upset by Item (2). ∎
It is well known that if an arithmetic frame $L$ has a unit, then $L_{d}$ is max-bounded (see, e.g., [8, before Prop. 3.3]). The following example shows that $L_{d}$ being max-bounded is a strictly weaker condition.
Example 7.6.
Let $L=\wp(\mathbb{N})$. Then $L$ is an arithmetic frame, and $da=a$ for all $a\in L$ since $L$ is Boolean, so $L=L_{d}$. The maximal elements of $L_{d}$ are exactly the coatoms. Therefore, $L_{d}$ is max-bounded since it is atomic. However, $L_{d}$ does not contain a unit since the only dense element is $1$, which is not compact.
We end this section by describing the Priestley analogue of having a unit and its relation to compactness of $\min Y_{d}$.
Theorem 7.7.
Let $L$ be an arithmetic frame and $X$ its L-space. The following are equivalent.
  1. (1)
    There is a unit in $L$.
  2. (2)
    There is a cofinal $U\in{\sf ClopSUp}(X)$.
  3. (3)
    There is $U\in{\sf ClopSUp}(X)$ such that $Y_{d}\subseteq U$.
The previous conditions imply the following equivalent conditions.
  1. (4)
    $\mathord{\uparrow}\min Y_{d}$ is a Scott upset.
  2. (5)
    $\min Y_{d}$ is compact.
If in addition $N_{d}$ is $d$-initial, then all five conditions are equivalent.
Proof.
[…]
$\min\mathord{\uparrow}(U\cap N_{d})\subseteq Y\cap N_{d}=Y_{d}.$
We show that $\min\mathord{\uparrow}(U\cap N_{d})=\min Y_{d}$. If $y\in\min\mathord{\uparrow}(U\cap N_{d})$ and $y^{\prime}\in Y_{d}$ such that $y^{\prime}\leq y$, then $y^{\prime}\in U\cap N_{d}$ because $Y_{d}\subseteq U$, so $y=y^{\prime}$ since
$y\in\min\mathord{\uparrow}(U\cap N_{d})=\min(U\cap N_{d}).$
Hence, $\min\mathord{\uparrow}(U\cap N_{d})\subseteq\min Y_{d}$. Conversely, if $y\in\min Y_{d}$ and $x\in\mathord{\uparrow}(U\cap N_{d})$ with $x\leq y$, then there is $y^{\prime}\in\min\mathord{\uparrow}(U\cap N_{d})\subseteq Y_{d}$ with $y^{\prime}\leq x\leq y$, so $y=y^{\prime}$ since $y\in\min Y_{d}$. Consequently, $\min Y_{d}=\min\mathord{\uparrow}(U\cap N_{d})$. Thus, $\mathord{\uparrow}\min Y_{d}=\mathord{\uparrow}(U\cap N_{d})$, and hence $\mathord{\uparrow}\min Y_{d}$ is a Scott upset.
[…]
Finally, suppose that $N_{d}$ is $d$-initial.
[…]
Remark 7.8.
In Example 8.11 we will show that the assumption in Theorem 7.7 that $N_{d}$ is $d$-initial is necessary. In fact, we will see that $\min Y_{d}$ may be compact Hausdorff without $L$ having a unit.
8. Hausdorffness of the maximal $d$-spectrum
In this final section, we give an example of an arithmetic frame $L$ with a unit such that $\max L_{d}$ is not Hausdorff, thus answering the question of [8] in the negative. In addition, we characterize exactly when $\max L_{d}$ is Hausdorff. Our characterization doesn’t require that $L$ has a unit, only that $\max L_{d}$ is locally compact. Under this assumption, we show that $\max L_{d}$ is Hausdorff iff it is stably locally compact, a condition that plays an important role in domain theory (see [21, Sec. VI.6]).
Example 8.1.
Consider the Stone-Čech compactification of the natural numbers $\beta\mathbb{N}$. Partition the natural numbers in countably many countable subsets $\mathbb{N}=X_{0}\cup X_{1}\cup X_{2}\cup\dots$, where $X_{i}=\{x_{i,0},x_{i,1},x_{i,2},\dots\}$. Then for each $X_{i}$, $\operatorname{cl}X_{i}$ is a clopen set of $\beta\mathbb{N}$ homeomorphic to $\beta\mathbb{N}$ (see, e.g., [18, p. 174]). Let $X_{i}^{*}=\operatorname{cl}(X_{i})\cap\mathbb{N}^{*}$ and let $Y_{\omega}=\{y_{0},y_{1},y_{2},\dots\}\cup\{\omega\}$ be the one-point compactification of a copy of the natural numbers. Consider now the disjoint union $X=\beta\mathbb{N}\sqcup Y_{\omega}$ and the partial order in Figure 2, where $X_{\omega}^{*}=\mathbb{N}^{*}\setminus\bigcup_{n\in\mathbb{N}}X_{n}^{*}$.
[…]
Figure 2. The Priestley space of an arithmetic frame whose maximal $d$-spectrum is not Hausdorff.
Our goal is to show that $X$ is an arithmetic L-space such that $\min Y_{d}$ is not Hausdorff. We have several things to verify.
Claim 8.2.
$X$ is a Priestley space.
Proof.
$X$ is a Stone space since it is the disjoint union of two Stone spaces. It remains to be shown that $X$ satisfies the Priestley separation axiom. For $x\in\beta\mathbb{N}$ and $x^{\prime}\in X$ with $x\not\leq x^{\prime}$, finding a clopen upset containing $x$ and missing $x^{\prime}$ is easy since $\beta\mathbb{N}$ is a clopen upset of $X$.
Let $y\in Y_{\omega}$ and $x\in X$ with $y\not\leq x$. Then $x\notin\mathord{\uparrow}\omega$, so $x\in\mathord{\downarrow}\operatorname{cl}X_{i}$ for some $i$ such that $y\neq y_{i}$. Consider $U=X\setminus\mathord{\downarrow}\operatorname{cl}X_{i}$. Since $\operatorname{cl}X_{i}$ is clopen, so is $\mathord{\downarrow}\operatorname{cl}X_{i}$. Thus, $U$ is a clopen upset separating $y$ from $x$. ∎
Claim 8.3.
$X$ is an L-space.
Proof.
It is sufficient to show that $\operatorname{cl}U$ is a clopen upset for each open upset $U\subseteq X$. Let $U$ be an open upset. Then $U\cap Y_{\omega}$ is open, and it is either empty or it must contain $\omega$. In both cases, $\operatorname{cl}(U)\cap Y_{\omega}=U\cap Y_{\omega}$, so $\operatorname{cl}U=U\cup\operatorname{cl}(U\cap\beta\mathbb{N})$, which is clearly a clopen upset. ∎
Claim 8.4.
The localic part of $X$ is $Y=Y_{\omega}\cup\mathbb{N}$.
Proof.
We have $Y\cap\mathbb{N}^{*}=\varnothing$ since $\mathord{\downarrow}x\cap\beta\mathbb{N}=\{x\}$ is not open for all $x\in\mathbb{N}^{*}$, so $Y\subseteq Y_{\omega}\cup\mathbb{N}$. For the converse, if $y\in Y_{\omega}\setminus\{\omega\}$ then $\mathord{\downarrow}y=\{y\}$ is clopen. Also, $\mathord{\downarrow}\omega=Y_{\omega}$ is clopen, so $Y_{\omega}\subseteq Y$. For $x\in X_{i}$, we have $\mathord{\downarrow}x=\{x,y_{i}\}$ is clopen. ∎
Claim 8.5.
Let $U\subseteq X$ be an upset. Then $U\in{\sf ClopSUp}(X)$ iff one of the following two conditions holds.
  1. (1)
    $U$ is a finite subset of $\mathbb{N}$.
  2. (2)
    $U\cap Y_{\omega}$ is cofinite, and $y_{i}\notin U$ implies $\operatorname{cl}(X_{i})\cap U$ is a finite subset of $X_{i}$.
Proof.
($\Rightarrow$) Suppose $U\in{\sf ClopSUp}(X)$ and $U\cap Y_{\omega}$ is not cofinite. Since $U\cap Y_{\omega}$ is clopen and not cofinite, $\omega\not\in U$. Hence, $U\cap Y_{\omega}=\varnothing$ since it is an upset. Therefore, $U\subseteq\beta\mathbb{N}=\operatorname{cl}\mathbb{N}$. By Item (1), $U\subseteq\mathbb{N}$, and by compactness $U$ is finite. Suppose now that $U\cap Y_{\omega}$ is cofinite. If $y_{i}\notin U$, then $X_{i}^{*}\cap U=\varnothing$, since otherwise $U$ can’t be Scott upset because $\min U\not\subseteq Y$ (see Claim 8.3). Therefore, $U\cap\operatorname{cl}(X_{i})\subseteq X_{i}$, and it has to be finite since it is compact.
($\Leftarrow$) If $U$ is a finite subset of $\mathbb{N}$, then $U\in{\sf ClopUp}(X)$ and $U\subseteq Y$ by Claim 8.4, so $U\in{\sf ClopSUp}(X)$. Now suppose (2) holds. Then $\omega\in U$ and $y_{i}\notin U$ implies $X_{i}^{*}\cap U=\varnothing$. Hence, $\min U\subseteq Y_{\omega}\cup\mathbb{N}=Y$, so it suffices to show that $U$ is clopen. Since $U\cap Y_{\omega}$ is cofinite it is clopen in $Y_{\omega}$. Moreover,
$U\cap\beta\mathbb{N}=\bigcup\{\operatorname{cl}(X_{i})\cap U\mid y_{i}\not\in U\}\cup\bigcup\{\operatorname{cl}(X_{i})\mid y_{i}\in U\}\cup X_{\omega}^{*}.$
By (2), $\bigcup\{\operatorname{cl}(X_{i})\cap U\mid y_{i}\not\in U\}=\bigcup\{X_{i}\cap U\mid y_{i}\not\in U\}\subseteq\mathbb{N}$ is finite and hence clopen. Also, $\beta\mathbb{N}\setminus(\bigcup\{\operatorname{cl}(X_{i})\mid y_{i}\in U\}\cup X_{\omega}^{*})$ is clopen since only finitely many $y_{i}\notin U$. Therefore, $\bigcup\{\operatorname{cl}(X_{i})\mid y_{i}\in U\}\cup X_{\omega}^{*}$ is clopen. Thus, $U\cap\beta\mathbb{N}$ is clopen, and so $U=(U\cap Y_{\omega})\cup(U\cap\beta\mathbb{N})$ is clopen. ∎
Claim 8.6.
$X$ is an algebraic L-space.
Proof.
Suppose $U\in{\sf ClopUp}(X)$. We need to show $U\subseteq\operatorname{cl}(\operatorname{core}U)$, so suppose $x\in U$. We consider three cases.
  1. (1)
    If $x\in Y_{\omega}$ then $\omega\in\mathord{\uparrow}x\subseteq U$, so $U\cap Y_{\omega}$ is cofinite. Then
    $x\in\mathord{\uparrow}(U\cap Y_{\omega})\in{\sf ClopSUp}(X)$
    by Item (2), and $\mathord{\uparrow}(U\cap Y_{\omega})\subseteq\operatorname{core}U$.
  2. (2)
    If $x\in\mathbb{N}$, then $\mathord{\uparrow}x=\{x\}\in{\sf ClopSUp}(X)$ by Item (1), and $\mathord{\uparrow}x\subseteq\operatorname{core}U$}.
  3. (3)
    Suppose $x\in\mathbb{N}^{*}$. Since $U$ is clopen in $X$, $U\cap\beta\mathbb{N}$ is clopen, and therefore
    $\operatorname{cl}(U\cap\mathbb{N})=U\cap\beta\mathbb{N}.$
    Hence, $x\in\operatorname{cl}(U\cap\mathbb{N})$. Suppose now that $V$ is a clopen neighborhood of $x$. Then $V\cap(U\cap\mathbb{N})\neq\varnothing$, but $U\cap\mathbb{N}=\bigcup\{\{n\}\mid n\in\mathbb{N}\cap U\}\subseteq\operatorname{core}U$ since $\{n\}\in{\sf ClopSUp}(X)$ by Item (1). Therefore, $x\in\operatorname{cl}\operatorname{core}U$. ∎
Claim 8.7.
$X$ is an arithmetic L-space.
Proof.
It suffices to show that $U\cap V\in{\sf ClopSUp}(X)$ for $U,V\in{\sf ClopSUp}(X)$ (see Item 2.15(1)), so suppose $U,V\in{\sf ClopSUp}(X)$. Then $U$ and $V$ satisfy one of the two conditions of Claim 8.5. If either $U$ or $V$ is a finite subset of $\mathbb{N}$, then so is their intersection. Suppose $U$ and $V$ both satisfy Item (2). Since a finite intersection of cofinite sets is cofinite, $U\cap V\cap Y_{\omega}$ is cofinite. If $y_{i}\notin U\cap V$, then either $y_{i}\notin U$ or $y_{i}\notin V$. Without loss of generality we may assume the former. Then $\operatorname{cl}(X_{i})\cap U\cap V\subseteq\operatorname{cl}(X_{i})\cap U$ is a finite subset of $X_{i}$. Thus, Item (2) holds for $U\cap V$, and so $U\cap V\in{\sf ClopSUp}(X)$. ∎
Claim 8.8.
$\min Y_{d}=Y_{\omega}\setminus\{\omega\}$.
Proof.
Observe that for each $y\in Y=\mathbb{N}\cup Y_{\omega}$, there is $x\in\max X$ with $\{y\}=\max(\mathord{\downarrow}x\cap Y)$. Therefore, $Y_{d}=\mathbb{N}\cup Y_{\omega}$ by Item (4). Consequently, $\min Y_{d}=Y_{\omega}\setminus\{\omega\}$. ∎
Claim 8.9.
$\min Y_{d}$ is not Hausdorff.
Proof.
${\sf ClopSUp}(X)$ forms a basis of $\min Y_{d}$ because $U\cap Y=\operatorname{core}U\cap Y$ for each $U\in{\sf ClopUp}(X)$. Consequently, it follows from Item (2) that $\min Y_{d}$ is equipped with the cofinite topology, which is not Hausdorff since $\min Y_{d}$ is infinite. ∎
Claims 8.7 and 8.9 yield the following:
Theorem 8.10.
There exist arithmetic L-spaces $X$ such that $\min Y_{d}$ is not Hausdorff.
As promised at the end of Section 7, we now demonstrate that $\min Y_{d}$ can be compact Hausdorff without $L$ having a unit.
Example 8.11.
Redefine the order in the space $X$ of Example 8.1 as in Figure 3.
[…]
Figure 3. The Priestley space of an arithmetic frame without a unit whose maximal $d$-spectrum is compact Hausdorff
A similar reasoning to the above yields that $X$ is an arithmetic L-space and its localic part is $Y\coloneqq\mathbb{N}\cup(Y_{\omega}\setminus\{\omega\})$. Observe that $Y_{d}=Y$ and $\min Y_{d}=\varnothing$, so it is trivially compact Hausdorff. However, $X$ has no cofinal clopen Scott upset since $X_{\omega}^{*}\subseteq\max X$ and $\mathord{\downarrow}X_{\omega}^{*}\cap Y=\varnothing$. Consequently, $X$ is the L-space of an arithmetic frame without units.
To characterize when $\min Y_{d}$ is Hausdorff, we recall (see, e.g., [21, Def. VI-6.7]) that a topological space $X$ is coherent if the binary intersection of compact saturated sets is compact. The space $X$ is stably locally compact if it is sober, locally compact, and coherent. A stably locally compact space is stably compact if it is compact, and spectral if in addition compact open sets form a basis. It is well known (see, e.g., [25, p. 75]) that a spectral space is Hausdorff iff it is $T_{1}$. The next lemma generalizes this result to stably locally compact spaces.
Lemma 8.12.
If $X$ is stably locally compact, then $X$ is Hausdorff iff $X$ is $T_{1}$.
Proof.
We only need to show the right-to-left implication. Suppose $x,y\in X$ are distinct. Let $\mathcal{K}_{x}=\{K\subseteq X\mid K\mbox{ is a compact saturated neighborhood of }x\}$ and define $\mathcal{K}_{y}$ similarly. It suffices to show that there exist $K_{x}\in\mathcal{K}_{x}$ and $K_{y}\in\mathcal{K}_{y}$ such that $K_{x}\cap K_{y}=\varnothing$. Since $X$ is $T_{1}$ and locally compact, for each $z\in X$ distinct from $x$, there is a compact saturated neighborhood $K$ of $x$ missing $z$. Therefore, $\bigcap\mathcal{K}_{x}=\{x\}$, and similarly $\bigcap\mathcal{K}_{y}=\{y\}$}. Consequently, $\bigcap\mathcal{K}_{x}\cap\bigcap\mathcal{K}_{y}=\varnothing$. By [21, Lem. VI-6.4], there exists a finite $\mathcal{K}\subseteq\mathcal{K}_{x}\cup\mathcal{K}_{y}$ such that $\bigcap\mathcal{K}=\varnothing$. Since $X$ is stably locally compact, $\mathcal{K}_{x}$ and $\mathcal{K}_{y}$ are directed. Therefore, there are $K_{x}\in\mathcal{K}_{x}$ and $K_{y}\in\mathcal{K}_{y}$ such that $K_{x}\cap K_{y}=\varnothing$. Thus, $X$ is Hausdorff. ∎
Theorem 8.13.
Let $X$ be an arithmetic L-space such that $\min Y_{d}$ is locally compact. Then $\min Y_{d}$ is Hausdorff iff $\min Y_{d}$ is stably locally compact.
Proof.
First suppose that $X$ is Hausdorff. Then $X$ is sober (see, e.g., [25, p. 43]). Let $K,M\subseteq X$ be compact saturated. Since $X$ is Hausdorff, $K,M$ are closed, so $K\cap M$ is closed. Since it is a closed subset of a compact set, it must be compact. Thus, $X$ is stably locally compact. Conversely, since $\min Y_{d}$ is $T_{1}$ by Proposition 6.7, $\min Y_{d}$ is Hausdorff by Lemma 8.12. ∎
Corollary 8.14.
Let $L$ be an arithmetic frame with a unit and $X$ its L-space.
  1. (1)
    $\min Y_{d}$ is Hausdorff iff $\min Y_{d}$ is stably locally compact.
  2. (2)
    $\max L_{d}$ is Hausdorff iff $\max L_{d}$ is stably locally compact.
Proof.
We only prove (1) as (2) follows from (1) and Theorem 6.6. Since $L$ has a unit, $\min Y_{d}$ is compact by Theorem 7.7. First suppose that $\min Y_{d}$ is Hausdorff. Then $\min Y_{d}$ is compact Hausdorff, and hence $\min Y_{d}$ is stably locally compact. Conversely, if $\min Y_{d}$ is stably locally compact, then Theorem 8.13 applies, and hence $\min Y_{d}$ is Hausdorff. ∎
We conclude the paper with several interesting open problems:
  • It remains open whether Theorem 8.13 can be reformulated as an equivalence between sobriety and Hausdorffness. Note that in Example 8.1, $\min Y_{d}$ is locally compact and coherent, but it fails to be Hausdorff solely because it is not sober.
  • It also remains open whether $\min Y_{d}$ is always locally compact (and/or coherent). In the absence of sobriety, local compactness of $\min Y_{d}$ is not equivalent to $\Omega(\min Y_{d})$ being a continuous frame (see, e.g., [25, p. 310]). This disparity emphasizes the importance of sobriety in these considerations. Indeed, as noted above, it is plausible that in this setting sobriety alone implies Hausdorffness.
  • Resolving the above questions requires developing a general method for identifying which topological spaces can be realized as $\min Y_{d}$. While each Stone space can be realized as such, it remains open whether the same can be said about each compact Hausdorff space (we note that it follows from [23] that each compact Hausdorff quasi F-space can be realized this way).